OpenStudy (anonymous):

x - y = 6 x = y + 2 Solve the system of equations by substitution. Answers: (4, -2) (6, 0) no solution coincident got confused on this -_- math is not my thing right now.

OpenStudy (solomonzelman):

\(\large\color{slate}{ x - y = 6 }\) \(\large\color{slate}{ x = \color{red}{ y + 2} }\) substitution into the first equation. \(\large\color{slate}{ (\color{red}{ y + 2}) - y = 6 }\)

OpenStudy (anonymous):

i did that and i got y - y = 8

OpenStudy (solomonzelman):

y-y=0, you know that, right?

OpenStudy (anonymous):

well excuse me Dx i am really bad with math

OpenStudy (solomonzelman):

but, you know that \(\large\color{slate}{ y-y=0 }\) . (just like, \(\large\color{slate}{ 1-1=0 }\) \(\large\color{slate}{ 4-4=0 }\) \(\large\color{slate}{ 72-72=0 }\) and etc.,) correct?

OpenStudy (anonymous):

yes... sorry i just get confused with algebra :\

OpenStudy (solomonzelman):

sure, no need to apologize. you have: \(\large\color{slate}{ y+2-y= 8 }\) right after the substitution. the y's cancel: \(\large\color{slate}{ 2= 8 }\)

OpenStudy (solomonzelman):

that means that there are no solutions.

OpenStudy (solomonzelman):

questions?

OpenStudy (anonymous):

oh ok thanks i was busting my head with that -_- ugh math

OpenStudy (anonymous):

but thank you so much

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

so would this also be no solutions? i am just checking this one before i redo my assignment 10x - 10y = 1 x = y - 3